OpenStudy (kramse18):

e^(-y)dx-(2y+xe^(-y))dy=0

OpenStudy (irishboy123):

\(e^{-y}dx-(2y+xe^{-y})dy=0\) doesn't look exact

OpenStudy (kramse18):

thats the problem

OpenStudy (loser66):

Why not? \(M(x,y) = e^{-y}\\M_y= -e^{-y}\) \(N(x,y) = -2y-xe^{-y}\\N_x=-e^{-y}\) Then, \(M_y=N_x\), it is exact.

OpenStudy (irishboy123):

cos its \(d\phi = 0 = \phi_x dx + \phi_y dy \) and with \(e^{-y}dx-(2y+xe^{-y})dy=0\) we have \(\phi_x = e^{-y}, \phi_{xy} = -e^{-y}\) \(\phi_y = -(2y+xe^{-y}), \phi_{yx} = -e^{-y}\) so it is exact

OpenStudy (kramse18):

yeah i got that. I'm stuck after that

OpenStudy (irishboy123):

\(\phi_x= e^{-y} \implies \phi = x e^{-y} + f(y)\) \(\phi_y = -2y-xe^{-y} \implies \phi = -y^2 + x e^{-y} + g(x)= x e^{-y} -y^2 + g(x)\) \(\phi = x e^{-y} -y^2 + C(x,y)\)

OpenStudy (irishboy123):

ask @Loser66 if that is right